OpenStudy (anonymous):

@satellite73 @Destinymasha @myininaya @esshotwired A recipe that makes 7 servings calls for 1/2 cup of apple juice and 2/3 cup of cranberry juice. How many cups of juice are there per serving? A.1/6 B.3/5 C.1 and 1/16 D. 8 and 1/6

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (anonymous):

help

OpenStudy (anonymous):

one half plus two thirds divided by 7 rather small i think

OpenStudy (anonymous):

what is \(\frac{1}{2}+\frac{2}{3}\) ?

OpenStudy (anonymous):

7/6

OpenStudy (anonymous):

right good and what is \(\frac{7}{6}\div 7\) ?

OpenStudy (anonymous):

1/6

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thanks @satellite73